OpenStudy (anonymous):

 In the formula A(t) = A0ekt , A is the amount of radioactive material remaining from an initial amount A0 at a given time t, and k is a negative constant determined by the nature of the material. A certain radioactive isotope decays at a rate of 0.125% annually. Determine the half-life of this isotope, to the nearest year

OpenStudy (anonymous):

A = A0*2^(t/T) A/A0 = 1 - 0.00125 = 0.99875 2^(-1/T) = 0.99875 -1/T = ln(0.99875) T = -1/ln(0.99875) T = 799.499 years